# Comment on “Properties and dynamics of generalized squeezed states”

### v1

@@ II Numerical Analysis
Here, $\hat{a}$ and $\hat{a}^{\dagger}$ are the photon annihilation and creation operators, respectively, so that $[\hat{a},\hat{a}^{\dagger}]=1$, and $r$ corresponds to the high-order squeezing parameter. Although in general $r\in\mathbb{C}$, for the analysis of average photon number $\langle\hat{a}^{\dagger}\hat{a}\rangle$, one can restrict without loss of generality to non-negative real values, $r\geq 0$. It is easy to verify that taking $n=1$, we recover the displacement operator expression, while considering $n=2$, we obtain a general expression of the two-photon squeezing operator. Our numerical analysis will focus on cases $n>2$, specifically for $n=3$ and $n=4$, as also considered in Ref. [1].
@@
For the cases $n=3$ and $n=4$, we refer to the states $\left|\textstyle{r_{3}}\right\rangle$ and $\left|\textstyle{r_{4}}\right\rangle$ as tri-squeezed and quadri-squeezed states, respectively. Now, we numerically calculate the average photon number of these states, as a function of the squeezing parameter $r$, for different truncation $N$ of the Fock basis. The results for both cases are shown in Figs. 1(a) $(n=3)$ and 1(b) $(n=4)$, and are consistent with those presented in Ref. [1]. For sufficiently small values of $r$, the average photon number for both cases does not change with the truncation. As $r$ grows, however, the curves for similar truncations match albeit their slope increases with $N$ without appearing to converge to any finite value. This in agreement with [3] where it was shown that after a finite value of $r$, the average photon number diverges. For even higher values of the squeezing parameter, the average photon number begins to oscillate, showing significantly different curves for closely spaced values of $N$. These behaviours, contrary to what is stated in Ref. [1], strongly suggests that the results in these regions are numerical artefacts lacking physical significance. In the next section we will test this hypothesis by performing an analytical analysis of these functions, estimating the radius of convergence and proving that for the physical regime, i.e. for $r$ such that $\langle\hat{a}^{\dagger}\hat{a}\rangle_{n}<\infty$, the average photon number is a non-decreasing function of $r$.
Figure 1: Behaviour of the average photon number $\langle\hat{a}^{\dagger}\hat{a}\rangle_{n}$ as a function of the squeezing parameter $r$. Panel (a) and (b) corresponds to the tri-squeezed, $\left|\textstyle{r_{3}}\right\rangle=\hat{U}_{3}(r)\left|\textstyle{0}\right\rangle$, and quadri-squeezed, $\left|\textstyle{r_{4}}\right\rangle=\hat{U}_{4}(r)\left|\textstyle{0}\right\rangle$, respectively. Different curves have been obtained under different truncation $N$ of the Fock basis, as indicated in the labels. Note the significant change of $\langle\hat{a}^{\dagger}\hat{a}\rangle_{n}$ for similar $N$ that suggests lack of convergence.
@@ III Convergence and monotonic analysis
Figure 2: (a) First $M=20$ and $M=10$ non-zero coefficients $c_{m}^{(3)}$ (orange dots) and $c_{m}^{(4)}$ (green dots) for $n=3$ and $n=4$, respectively, as a function of $m$. An exponential fit for the last five points for each case have been plotted as straight lines. (b) Average photon number for the tri-squeezed state as a function of the squeezing parameter for different approximations. Orange and green lines correspond to a numerical computation using $N=60000$ and $N=60001$ Fock basis, respectively. The dashed black line represents the Taylor series considering the first $M=20$ terms. The dashed red line denotes the estimated radius of convergence $R_{3}$ for the Taylor series.
@@
where $[\hat{A},\hat{B}]_{m+1}=[A,[A,B]_{m}]$ is the nested commutator of order $m$ and $[A,B]_{m=0}=B$. This allows us to write the average photon number as infinite series,
@@
where the coefficients $c_{m}^{(n)}$ result from $c_{m}^{(n)}=\frac{1}{m!}\left\langle\textstyle{0}\right|[(\hat{a}^{\dagger,n}-\hat{a}^{n}),\hat{a}^{\dagger}\hat{a}]_{m}\left|\textstyle{0}\right\rangle$. For $n\geq 3$, the higher-than-quadratic boson operators prevents a closed expression for Eq. (5). Yet, one can compute the coefficients $c_{m}^{(n)}$ up to some order $M$. In Fig. 2(a), we present the first $M=20$ and $M=10$ non-zero coefficients $c_{m}^{(3)}$ and $c_{m}^{(4)}$ for $n=3$ and $n=4$, respectively, as a function of $m$. Note that, by symmetry, the odd coefficients are zero $c_{2m+1}^{(n)}=0$. The coefficients have been computed using the Mathematica-based software DiracQ [5]. As it can observed in Fig. 2(a), the coefficients grow exponentially with $m$, i.e. $c_{m}^{(n)}\propto e^{\alpha_{n}m}$. A fit, represented as straight lines in the graph, allows us to obtain the coefficients $\alpha_{n}$, resulting in $\alpha_{3}=1.95(2)$ and $\alpha_{4}=3.4(1)$. Therefore, we can estimate the radius of convergence of the average photon number as $R^{-1}_{n}=\lim_{m\rightarrow\infty}|c_{m}^{(n)}|^{1/m}=e^{\alpha_{n}}$, which results in $R_{3}\approx 0.14$ and $R_{4}\approx 0.03$. In Fig. 2(b) we have represented the average photon number for the tri-squeezed state, using the numerical approximation employing a truncated space, and the Taylor expansion keeping the first $M=20$ terms (cf. Eq. (5)). The radius of convergence $R_{3}$ is also shown as a vertical line. For values $r>R_{3}$, the state contains an infinite number of photons, and therefore it becomes non-physical [3]. Due to the divergence for $r>R_{3}$, results based on an artificially truncated Fock basis or truncated series are misleading numerical artefacts that lack physical meaning. This unphysical regime indicates that the simple interaction assumed to generate the gate $\hat{U}_{3}(r)$ breaks down, as for example requiring to explicitly include the quantized pump field in a $3$-photon downconversion as discussed in Ref. [3]. We note that the oscillations take place in this regime. For $r<R_{3}$, however, these different approximations to the average photon number match, as it can be seen in Fig. 2(b), showing a growing number of photons for increasing $r$. This corresponds to the physical region. Although not explicitly shown, the situation for $n=4$ is equivalent.
@@
$\hat{A}_{n}=\sum_{k=1}^{n}k!C_{n,k}^{2}\prod_{j=0}^{n-k-1}(\hat{a}^{\dagger}\hat{a}-j),\ {\rm for}\ n\geq 1,$ (7)
where $C_{n,k}$ represents the binomial coefficient, $C_{n,k}=\frac{n!}{k!(n-k)!}$. Since $(\hat{a}^{\dagger}\hat{a})^{k}$ with $k\in\mathbb{N}$ is non-negative, the operator $\hat{A}_{n}$ is strictly positive, $\hat{A}_{n}>0,\ \forall n\in\mathbb{N}^{+}$. In particular, for $n=1,2,3$ and $4$, this operator takes the form
$\hat{A}_{1}$ $=1,$ (8)
$\hat{A}_{2}$ $=4\hat{a}^{\dagger}\hat{a}+2,$ (9)
$\hat{A}_{3}$ $=9(\hat{a}^{\dagger}\hat{a})^{2}+9\hat{a}^{\dagger}\hat{a}+6,$ (10)
$\hat{A}_{4}$ $=16(\hat{a}^{\dagger}\hat{a})^{3}+24(\hat{a}^{\dagger}\hat{a})^{2}+56\hat{a}^{\dagger}\hat{a}+24.$ (11)
In this manner, we arrive to
@@
Assume $\langle\hat{a}^{\dagger}\hat{a}\rangle_{n}$ to be a non-monotonic function with $|r|$. This means that $\langle\hat{a}^{\dagger}\hat{a}\rangle_{n}$ decreases after some value $|\tilde{r}|>0$, reaching a local maximum at $|\tilde{r}|$. However, $\frac{d^{2}}{d|r|^{2}}\langle\hat{a}^{\dagger}\hat{a}\rangle_{n}>0\ \forall|r|$, and so also at $|\tilde{r}|$. This rules out the existence of local maxima, therefore contradicts the first assumption, i.e. $\langle\hat{a}^{\dagger}\hat{a}\rangle_{n}$ is a monotonic function. Moreover, since $\langle\hat{a}^{\dagger}\hat{a}\rangle_{n}\geq 0,\forall r\in\mathbb{C},\forall n\in\mathbb{N}^{+}$, being zero for $|r|=0$, implies that $\langle\hat{a}^{\dagger}\hat{a}\rangle_{n}$ is a non-decreasing function with $|r|$. ∎

### v2

@@ II Numerical Analysis
Here, $\hat{a}$ and $\hat{a}^{\dagger}$ are the photon annihilation and creation operators, respectively, so that $\hat{a},\hat{a}^{\dagger}]=1$, and $r$ corresponds to the high-order squeezing parameter. Although in general $r\in\mathbb{C}$, for the analysis of average photon number $\langle\hat{a}^{\dagger}\hat{a}\rangle$, one can restrict without loss of generality to non-negative real values, $r\geq 0$. It is easy to verify that taking $n=1$, we recover the displacement operator expression, while considering $n=2$, we obtain a general expression of the two-photon squeezing operator. Our numerical analysis will focus on cases $n>2$, specifically for $n=3$ and $n=4$, as also considered in Ref. [1].
@@
For the cases $n=3$ and $n=4$, we refer to the states $\left|\textstyle{r_{3}}\right\rangle$ and $\left|\textstyle{r_{4}}\right\rangle$ as tri-squeezed and quadri-squeezed states, respectively. Now, we numerically calculate the average photon number of these states, as a function of the squeezing parameter $r$, for different truncation $N$ of the Fock basis. The results for both cases are shown in Figs. 1(a) $(n=3)$ and 1(b) $(n=4)$, and are consistent with those presented in Ref. [1]. For sufficiently small values of $r$, the average photon number for both cases does not change with the truncation. As $r$ grows, however, the curves for similar truncations match albeit their slope increases with $N$ without appearing to converge to any finite value. This in agreement with [3] where it was shown that after a finite value of $r$, the average photon number diverges. For even higher values of the squeezing parameter, the average photon number begins to oscillate, showing significantly different curves for closely spaced values of $N$. These distinct shapes depend on the parity of $N^{\prime}=\lfloor(N-1)/n\rfloor$ with $\lfloor x\rfloor=\max\{n\in\mathbb{N}:n\leq x\}$, revealing how the enlargement of the Hilbert space affects the system when the newly added Fock state can be populated, i.e. when $(N-1)/n\in\mathbb{N}$. These behaviours, contrary to what is stated in Ref. [1], strongly suggests that the results in these regions are numerical artefacts lacking physical significance. This includes the average photon number $\langle\hat{a}^{\dagger}\hat{a}\rangle_{n}$, as well as the Husimi and Wigner functions. In the next section we will test this hypothesis by performing an analytical analysis of these functions, estimating the radius of convergence and proving that for the physical regime, i.e. for $r$ such that $\langle\hat{a}^{\dagger}\hat{a}\rangle_{n}<\infty$, the average photon number is a non-decreasing function of $r$.
Figure 1: Behaviour of the average photon number $\langle\hat{a}^{\dagger}\hat{a}\rangle_{n}$ as a function of the squeezing parameter $r$. Panels (a) and (b) corresponds to the tri-squeezed, $\left|\textstyle{r_{3}}\right\rangle=\hat{U}_{3}(r)\left|\textstyle{0}\right\rangle$, and quadri-squeezed, $\left|\textstyle{r_{4}}\right\rangle=\hat{U}_{4}(r)\left|\textstyle{0}\right\rangle$, respectively. Different curves have been obtained under different truncation $N$ of the Fock basis, as indicated in the labels. Note the significant change of $\langle\hat{a}^{\dagger}\hat{a}\rangle_{n}$ for similar NN that suggests lack of convergence.
@@ III Convergence and monotonic analysis
Figure 2: (a) First $M=20$ and $M=10$ coefficients $c_{m}^{(3)}$ (orange dots) and $c_{m}^{(4)}$ (green dots) for $n=3$ and $n=4$, respectively, as a function of $m$. Note that, by symmetry, the odd coefficients are zero $c_{2m+1}^{(n)}=0$. An exponential fit for the last five points for each case have been plotted as straight lines. (b) Average photon number for the tri-squeezed state as a function of the squeezing parameter for different approximations. Orange and green lines correspond to a numerical computation using $N=60000$ and $N=60001$ Fock basis, respectively. The dashed black line represents the Taylor series considering the first $M=20$ terms. The dashed red line denotes the estimated radius of convergence $R_{3}$ for the Taylor series.
@@
where $\hat{A},\hat{B}]_{m+1}=[A,[A,B]_{m}]$ is the nested commutator of order $m$ and $A,B]_{m=0}=B$. This allows us to write the average photon number as infinite series,
@@
where the coefficients $c_{m}^{(n)}$ result from $c_{m}^{(n)}=\frac{1}{m!}\left\langle\textstyle{0}\right|[(\hat{a}^{\dagger,n}-\hat{a}^{n}),\hat{a}^{\dagger}\hat{a}]_{m}\left|\textstyle{0}\right\rangle$. For $n\geq 3$, the higher-than-quadratic boson operators prevents a closed expression for Eq. (5). Yet, one can compute the coefficients $c_{m}^{(n)}$ up to some order $M$. In Fig. 2(a), we present the first $M=20$ and $M=10$ coefficients $c_{m}^{(3)}$ and $c_{m}^{(4)}$ for $n=3$ and $n=4$, respectively, as a function of $m$. Note that, by symmetry, the odd coefficients are zero $c_{2m+1}^{(n)}=0$. The coefficients have been computed using the Mathematica-based software DiracQ [5]. As it can observed in Fig. 2(a), the coefficients grow exponentially with $m$, i.e. $c_{m}^{(n)}\propto e^{\alpha_{n}m}$. A fit, represented as straight lines in the graph, allows us to obtain the coefficients $\alpha_{n}$, resulting in $\alpha_{3}=1.95(2)$ and $\alpha_{4}=3.4(1)$. Therefore, we can estimate the radius of convergence of the average photon number as $R^{-1}_{n}=\lim_{m\rightarrow\infty}|c_{m}^{(n)}|^{1/m}=e^{\alpha_{n}}$, which results in $R_{3}\approx 0.14$ and $R_{4}\approx 0.03$. These approximated $R_{3,4}$ also apply to the results shown in Ref. [1], as we used the same definition for the generalized squeezing operator $\hat{U}_{n}(r)$. In Fig. 2(b) we have represented the average photon number for the tri-squeezed state, using the numerical approximation employing a truncated space, and the Taylor expansion keeping the first $M=20$ terms (cf. Eq. (5)). The radius of convergence $R_{3}$ is also shown as a vertical line. For values $r>R_{3}$, the state contains an infinite number of photons, and therefore it becomes non-physical [3]. Due to the divergence for $r>R_{n}$, results out of the convergence radius using artificially truncated Fock basis or series are misleading numerical artefacts that lack physical meaning, including the Husimi and Wigner functions shown in [1]. This unphysical regime indicates that the simple interaction assumed to generate the gate $\hat{U}_{3}(r)$ breaks down, as for example requiring to explicitly include the quantized pump field in a $3$-photon downconversion as discussed in Ref. [3]. We note that the oscillations take place in this regime. For $r<R_{3}$, however, these different approximations to the average photon number match, as it can be seen in Fig. 2(b), showing a growing number of photons for increasing $r$. This corresponds to the physical region. Although not explicitly shown, the situation for $n=4$ is equivalent.
@@
$\hat{A}_{n}=\sum_{k=1}^{n}k!C_{n,k}^{2}(\hat{a}^{\dagger,n-k}\hat{a}^{n-k}),\ {\rm for}\ n\geq 1,$ (7)
where $C_{n,k}$ represents the binomial coefficient, $C_{n,k}=\frac{n!}{k!(n-k)!}$. Since $(\hat{a}^{\dagger,n-k}\hat{a}^{n-k})$ with $n-k\in\mathbb{N}$ is non-negative, the operator $\hat{A}_{n}$ is strictly positive, $\hat{A}_{n}>0,\ \forall n\in\mathbb{N}^{+}$. Note that one can also see that $\hat{A}_{n}>0$ by expressing the commutator $\hat{A}_{n}$ in the Fock basis. In this manner, we arrive to
@@
[…]